OpenStudy (anonymous):

What is 34% of 50? A. 7 B. 17 C. 34 D. 68

OpenStudy (anonymous):

50*34/100=34/2=17%

OpenStudy (anonymous):

Forty-two is what percent of 70? A. 21% B. 60% C. 74% D. 167%

OpenStudy (anonymous):

17

OpenStudy (anonymous):

42/70 = 60%

OpenStudy (anonymous):

Forty-five is 30% of what number? A. 135 B. 67 C. 150 D. 45

OpenStudy (anonymous):

45/150 = 30% so 45 is 30% of 150

OpenStudy (anonymous):

Ben bought a digital camera for $90. The tax rate is 6%. Which proportion can be used to find the amount of tax Ben will pay?

OpenStudy (anonymous):

90*6% = 5.4 he will pay $5.40 for tax

OpenStudy (anonymous):

Lee had a total of 40 e-mails on Monday. Eight of the e-mails were from her friend Cara. What percent of the e-mails were from Cara? A. 8% B. 20% C. 32% D. 40%

OpenStudy (anonymous):

8/40 = 20%

OpenStudy (anonymous):

What is the percent of decrease from 320 to 32? A. 10% B. 90% C. 100% D. 288%

OpenStudy (anonymous):

poop no tis is too old and done whaa i would have helped

OpenStudy (anonymous):